If you use a 0.05 level of significance in a two-tail hypothesis test, what decision will you make if ZSTAT=−0.76? Please show work in excel with explanations. thank you Formulas as well
Answer #1

P value for this two tail hypothesis is 2* P(:< -0.76) = 2 * NORMSDI ST(-0.76) = 2 * 0.2236 = 0.4472 > o= 0,05

As P value is greater than alpha=0.05 we fail to reject the null hypothesis